OpenStudy (anonymous):

find dx/dy if X3 +x tan-1(y)=e y

myininaya (myininaya):

I can't read this very well.

myininaya (myininaya):

what do the sup's mean?

myininaya (myininaya):

do you mean x^3+arctan(y)=e^y?

OpenStudy (anonymous):

its mean quadrate, yes, please help its for my brother math homework

myininaya (myininaya):

Do you know power rule? And I guess we are looking at x as a function of y. \[(x^n)'=nx^{n-1} \frac{dx}{dy}\] \[(\arctan(x))'=\frac{1}{1+x^2} \cdot \frac{dx}{dy}\] \[(e^x)'=e^x \frac{dx}{dy} \] I hope that helps you and your brother get started.

OpenStudy (anonymous):

thank you so much, we will try get started